OpenStudy (anonymous):

if cos theta=-5/11 and theta is in quadrent 3 find cot(theta)tan(theta

OpenStudy (jdoe0001):

\(\bf cos(\theta)=\cfrac{-5}{11}\to \cfrac{adjacent}{hypotenuse}\to \cfrac{x}{radius}\qquad thus\) so just find the opposite side, or "b", using the pythagorean theorem keeping in mind that \(\bf cot(\theta)=\cfrac{x}{y}\to \cfrac{b}{a}\qquad tan(\theta)=\cfrac{y}{x}\to \cfrac{a}{b}\)

OpenStudy (jdoe0001):

hmm actually keeping in mind that \(\bf cot(\theta)=\cfrac{x}{y}\to \cfrac{a}{b}\qquad tan(\theta)=\cfrac{y}{x}\to \cfrac{b}{a}\)

OpenStudy (anonymous):

but it says that cot and tan are multiplied together. when doing the math i got that tan is the squareroot of 96/-5

OpenStudy (jdoe0001):

well ... if tangent and cotangent are multiplying by each other, using the same angle.. then they're always \(\bf cot(\theta)\cdot tan(\theta)\implies \cfrac{1}{\cancel{ tan(\theta) }}\cdot \cancel{ tan(\theta) }\) thus, I'm thinking you're asked to get each separately maybe

OpenStudy (anonymous):

does this help?

OpenStudy (anonymous):

i tried to cancel it out, but when i plugged it in it said that I was incorrect.